Calculating ROI – Use the following information for The Regal Company to compute ROI.

Net operating income ($50,000)

Average operating assets ($230,000)

Sales ($535,000)
Operating expenses($485,000)

Compute Regal's new ROI

1a. Assume that Regal's manager invests in a $30,000 piece of equipment that increases sales by $35,000, while increasing operating expenses by $15,000 and reports the following:

Net operating income ($50,000)

Average operating assets($230,000)

Sales($535,000)
Operating expenses ($485,000)

Compute Regal’s new ROI.

1b. Computing Residual Income - The Retail Division of Zephyr, Inc. has average operating assets of $100,000andisrequiredtoearnareturnof20%ontheseassets. Inthecurrentperiod,thedivision earns $30,000. Compute residual income for the Retail Division.

1c.The Wholesale Division of Zephyr, Inc. has average operating assets of $1,000,000 and is also required to earn a return of 20% on these assets. In the current period, the division earns $220,000. Compute residual income for the Wholesale Division.

1d. Which division is performing better?
Answer #1
1a) ROI = Net operating income/Average operating assets = 50000/230000 = 21.74%
Note: It is assumed that the figures of $50000 and
$230000 include the operating income of $20000
[35000-150000] and increase in operating assets of
$30000 for the new investment.
If it is extra, the new ROI = (50000+20000)/(230000+30000) = 26.92%
1b) RI for retail division = NOI-Average operating assets*Required return = 30000-100000*20% = $        10,000
1c) RI of wholesale division = 220000-1000000*20% = $        20,000
1d) Based on residual income, Wholesale division is performing
better.